# GrAFFORD: A Benchmark Dataset for Testing the Knowledge of Object Affordances of Language and Vision Models

### v1

@@ Abstract
We investigate the knowledge of object affordances in pre-trained language models (LMs) and pre-trained Vision-Language models (VLMs). Transformers-based large pre-trained language models (PTLM) learn contextual representation from massive amounts of unlabeled text and are shown to perform impressively in downstream NLU tasks. In parallel, a growing body of literature shows that PTLMs fail inconsistently and non-intuitively, showing a lack of reasoning and grounding. To take a first step toward quantifying the effect of grounding (or lack thereof), we curate a novel and comprehensive dataset of object affordances – GrAFFORD, characterized by 15 affordance classes. Unlike affordance datasets collected in vision and language domains, we annotate in-the-wild sentences with objects and affordances. Experimental results reveal that PTLMs exhibit limited reasoning abilities when it comes to uncommon object affordances. We also observe that pre-trained VLMs do not necessarily capture object affordances effectively. Through few-shot fine-tuning, we demonstrate improvement in affordance knowledge in PTLMs and VLMs. Our research contributes a novel dataset for language grounding tasks, and presents insights into LM capabilities, advancing the understanding of object affordances. Codes and data are available at https://github.com/sayantan11995/Affordance
@@ 1 Introduction
The task of object affordances has been well studied in Computer Vision and Robotics Zhu et al. (2014); Kjellström et al. (2011); Gupta et al. (2009). However, infusing such knowledge while learning textual representation is hard; as in NLP, we lack corresponding images (or videos) which may provide necessary shape, property information to predict affordances. The state-of-the-art transformers-based large pre-trained language models (PTLMs) learn textual representation from massive amounts of unlabeled text and are shown to perform impressively in downstream NLU tasks. In parallel, a growing body of literature shows that PTLMs fail inconsistently and non-intuitively, showing a lack of reasoning and grounding. In NLP, these results prompted authors in Bender and Koller (2020) to reiterate the gap between form and meaning.
The authors argue that language models which are exposed to only text (surface form) may never truly understand meaning, as PTLMs are unaware of true groundings of the surface text. Our goal is to first precisely quantify the gap, capturing one aspect of grounding, aka affordability. To this end, we take inspiration from vision and robotics research and explore affordance properties of entities or objects in text, which a model with proper grounding should be able to estimate. We propose a novel text-to-affordance dataset, and explore to what extent textual representations learnt by PTLMs and VLMs can enable reasoning with affordance of entities mentioned in text.
@@
As an example, for a sentence “an apple in the tree”, we should infer that the “apple” can be thrown, and is rollable. However we cannot roll an “apple logo”. In computer vision and robotics efforts, an accompanying image (or video) often provides necessary information about shape and physical properties of an entity, which can be used to predict affordances Zhu et al. (2014). However such information is absent in NLP tasks. To capture this nuance, we annotate crowdsourced text intended for other tasks (such as NLI) with the objects and affordances. We use 15 affordance classes from Zhu et al. (2014). Through extensive pilot studies, we train a set of annotators using the toloka.ai platform. We choose $25$ highly-skilled annotators who annotates a total of $2368$ sentence-object pairs with 15 affordance classes, on a 0-3 Likert-like scale. We name this novel dataset GrAFFORD.We use the dataset for zero-shot evaluations of small LMs, two open-source LLMs and also some VLMs. We evaluate the effect of few-shot fine-tuning on few PTLMs and VLM. Our contributions can be summarized as follows.
@@
  • We curate a novel large scale crowdsource based text to affordance dataset – GrAFFORD, consisting of 2368 sentence-object pairs and 15 affordance classes for each pair.
@@
  • We perform zero-shot evaluation of many state-of-the art PTLMs along with a few VLMs in different settings to identify the extent to which they gain the knowledge of affordance using our dataset. We further ensemble the VLM and the PTLM predictions to examine whether pre-training with images can enrich affordance prediction from text.
@@
  • We also fine-tune few PTLMs on a small subset of our data as well as on some commonsense reasoning tasks to understand how quickly the affordance knowledge get scaled up and how far the affordances are related to common sense knowledge. In addition, we examine the in-context learning (ICL) ability of few of the SOTA generative LLMs and VLMs in affordance prediction task.
  • We provide with a comprehensive analysis of the model prediction and explore how the linguistic ambiguity of the objects can affect the models’ ability of affordance prediction.
@@ 2 Related work
Natural language grounding. In vision and robotics, grounding has traditionally referred to as locating and identifying text expressions within images (or videos). Yu et al. (2015) introduce referring expression grounding which grounds referring expressions within given images via jointly learning the region visual feature and the semantics embedded in referring expressions. Chen et al. (2017) present phrase grounding which aims to locate referred targets by corresponding phrases in natural language queries. Shridhar and Hsu (2018) employ expressions generated by a captioning model Johnson et al. (2016), gestures, and a dialog system to ground targets. Another line of work Bastianelli et al. (2016); Alomari et al. (2017) explored non-dialog based methods to ground text queries.
Reasoning about object affordances. Sun et al. (2014) learnt object affordances through human demonstrations & Kim and Sukhatme (2014) deduced affordance through extracted geometric features from point cloud segments. Zhu et al. (2014) reasoned affordance through querying the visual attributes, physical attributes, and categorical characteristics of objects by employing a pre-built knowledge base. Myers et al. (2015) perceive affordance from local shape and geometry primitives of objects. Recent methods employed deep learning approaches to detect object affordance. Roy and Todorovic (2016) used a multi-scale CNN to extract mid-level visual features and combine them to segment affordances from RGB images. Sawatzky et al. (2017) regard affordance perception as semantic image segmentation and adopt a CNN based architecture to segment affordances from weakly labeled images. Nguyen et al. (2016); Mi et al. (2019) utilized features from a pre-trained CNN model to predict object affordances. Nguyen et al. (2017) applied an object detector, CNN and dense conditional random fields to detect object affordance from RGB images.
@@
Probing methods. Talmor et al. (2020) utilized probing and employed Multi-choice MLM (Masked Language Modelling) and Multi-choice QA (Question Answering) setup to capture reasoning capabilities of pre-trained Language Models. Yang et al. (2022) examine zero-shot prediction performances on different tasks by LLM through novel visual imagination. Aroca-Ouellette et al. (2021) highlighted the shortcomings of state-of-the-art pre-trained models in physical reasoning, with a further performance decline observed when incorporating option shuffling and superlatives in reasoning questions. Liu et al. (2022) proposed a novel spatial common sense probing framework to investigate object scales and positional relationship knowledge in text-based pre-trained models and models with visual signals.
3 GrAFFORD dataset construction
@@
Preprocessing. We select $20000$ sentences from a crowdsourced English dataset (XNLI English) Conneau et al. (2018) [^1] We choose XNLI as a source to facilitate multilingual extensions of our dataset. and extract the noun phrases using the Stanford CoreNLP tool. As we restrict to the affordances that humans can directly perform, we filter the phrases which do not represent a tangible object (using ConceptNet). We manually filter out objects that cannot be acted upon directly by humans (such as school, building). After this preprocessing, we obtain a set of sentence-object pairs ($\langle x_{i},o_{i}\rangle$), where the sentence acts as the context for the corresponding object. Each sentence on average has 2-3 such objects. We use the 15 predefined affordance classes from Zhu et al. (2014) to label each sentence-object pair for annotation.
Pilot studies & annotator training. We annotate the dataset using the Toloka platform [^4] https://toloka.ai/. We design an interface on this platform, which contained clear instructions and examples for annotating the data. We conduct two rounds of pilot studies to analyze the subjective understanding of the annotators and, thereby, filter out the high quality, serious annotators. For the first pilot study, we present the annotators with the smaller 62 sentence-object pairs and ask them to label the instance with each affordance class on a scale of 00 to $3$, indicating whether or not the affordance can be performed on the object. Here, 0-1 indicates that the affordance cannot be performed (high-low) and 2-3 indicates that the affordance can be performed low-high). We will further use these 62 synthetic sentence-object pairs for few-shot training. For quality control, we select the top 90% of the available annotators in the platform, who are proficient in English, and use computers to complete the tasks [^5] We exclude mobile-users as we believe the instructions may not appear clearly on mobile devices.. A total of 15 annotators labelled the data, and all of them were incentivized uniformly. After the first pilot, we find that there is an extremely poor agreement among the annotators, and the overall precision is around 28%. Therefore, we moved on to a second pilot study. Here, we use all the 62 sentence-object pairs from the previous study, along with 32 randomly selected sentence-object pairs from the XNLI data. We use the top 30% of the annotators (based on the quality determined by the platform) available on the platform, while other criteria remained the same. We annotate 32 sentence-object pairs ourselves, and use all the labelled examples as control data points to guide the annotators while labelling. A total of 114 annotators participated in this version of the pilot study. We assign a specific skill to the annotators who attained more than $30\%$ precision and $30\%$ recall. In total, 48 annotators passed this criteria. Through initial pilot studies, we learnt that without grounded images, the task appears quite subjective to annotators. The main goal of the pilot studies have been to understand the annotators’ quality, their comprehension of the task, and their preferences for incentives per task. We have also conducted two additional AMA (Ask Me Anything) sessions with interested annotators to further clarify the task.
Final annotation. In the final phase, we conduct the annotation on a larger set of sentence-object pairs, carefully selecting a total of 2,368 pairs. To ensure diverse perspectives and minimize bias, we engage 25 skilled annotators in this phase. Three annotators independently annotated each of the sentence-object pairs. Each annotator meticulously evaluated the affordance classes for every pair, contributing to a comprehensive annotation of the dataset. We perform the annotations in phases and complete the full task over 10 phases. We measure class-wise agreement and average agreement across all classes after each annotation phase to ensure the quality of the annotations. The overall statistics for this currently constructed dataset – GrAFFORD is in Tab. 1. Throughout the data processing pipeline, we put meticulous attention to the quality control, including the use of pilot studies, iterative annotation refinement, and manual filtering. These measures ensure that the dataset is comprehensive, accurate, and aligned with the objectives of the study. Overall, our GrAFFORD dataset consists of $2368$ sentence-object pairs having $\sim 100k$ annotations (2368 $\times$ 15 $\times$ 3).
@@
Table 1: GrAFFORD dataset statistics.
@@
GrAFFORD data exploration. We observe that classes such as ‘Grasp’, ‘Lift’, ‘Throw’, ‘Push’, and ‘Watch’ are the most common affordances for the objects present in the dataset (see Figure 1). Most frequent objects and their corresponding agreement scores are shown in Appendix A.7 Fig. 6. We observe, agreement scores are fairly uniform ($0.5$-$0.6$) for frequent objects, with high agreement for some frequent objects ($0.8$ for “the movie”). In Figure 7 (see Appendix A.6), we also see that ‘Grasp’, ‘Lift’, and ‘Throw’ are highly correlated classes. There is similar positive correlation between the class ‘SitOn’ and ‘Ride’, and some correlation between ‘Watch’ and ‘LookThrough’.
In Table 2, we list down the affordance classes based on the annotator agreement score, and divide it into three categories to understand which of the affordance classes pose the most and least difficulties for the human annotators. We observe that the classes - ‘Watch’, ‘SitOn’, and ‘TypeOn’ are the most difficult to disambiguate.
Agreement category
Affordance classes
Objects
Object-affordance pair
High agreement (>0.6)
Row, Feed, Ride, Fix
the horse, striped white shirts, a brown paper sack, Chinese lanterns, Adrin’s sword, The movie
breakfast-Feed, a horse-Watch, crops-Fix, sports-Grasp, sports-Lift, sports-Push, the phone-Feed, football-Ride
Medium agreement (>0.45 & <0.6)
Throw, PourFrom, WriteWith, LookThrough, Lift, Grasp, Play, Push
A red flag, An arrow, Art, Automatic weapons, Babies, Black-and-white TV
computers-WriteWith, cats-Feed, football-Play, book-WriteWith, the door-Push
Low agreement (<0.45)
Watch, SitOn, TypeOn
Brandy from Spain, stone circles, iron, batteries, his fist, historical artifact, gift, olive oil, outdoor tables, bumper sticker on a car
weapon-Push, The table-Lift, boat-Fix, paintings-LookThrough, cats-Throw
Table 2: Annotator agreement based on the difficulty in disambiguating different affordance classes and objects.

### v2

# Text2Afford: Probing Object Affordance Prediction abilities of Language Models solely from Text

@@ Abstract
We investigate the knowledge of object affordances in pre-trained language models (LMs) and pre-trained Vision-Language models (VLMs). A growing body of literature shows that PTLMs fail inconsistently and non-intuitively, demonstrating a lack of reasoning and grounding. To take a first step toward quantifying the effect of grounding (or lack thereof), we curate a novel and comprehensive dataset of object affordances – Text2Afford, characterized by 15 affordance classes. Unlike affordance datasets collected in vision and language domains, we annotate in-the-wild sentences with objects and affordances. Experimental results reveal that PTLMs exhibit limited reasoning abilities when it comes to uncommon object affordances. We also observe that pre-trained VLMs do not necessarily capture object affordances effectively. Through few-shot fine-tuning, we demonstrate improvement in affordance knowledge in PTLMs and VLMs. Our research contributes a novel dataset for language grounding tasks, and presents insights into LM capabilities, advancing the understanding of object affordances.  [^1] Code and Data are available anonymously at https://anonymous.4open.science/r/Text2Afford
@@ 1 Introduction
Gaining the knowledge of object affordances while learning textual representation from large corpora maybe hard; as in NLP, we lack corresponding images (or videos) which provides necessary visual cues such as shape, color, and texture to predict affordances. This lack of mapping or rather grounding ability has been noted by many researchers in the context of large pretrained language models (PTLMs). Authors in Bender and Koller (2020) have pointed the lack of symbol grounding to be a fundamental factor behind PTLMs failing to grasp meaning from form (surface form text). The authors argue that language models which are exposed to only text (surface form) may never truly understand meaning, as PTLMs are unaware of possible groundings of the surface text.
Figure 1: Overview of Text2Afford with its derived tasks.
Most current NLP datasets and tasks are not designed to evaluate grounding, as it is hard to evaluate grounding without any visual context. Here, we aim to quantify the ability of pretrained models to learn affordances – which in turn requires the ability to ground symbols in text to real-world objects. In other words, grounding ability from text can enable understanding and reasoning about the physical properties of an object, which may help predict affordances.
  Dataset Train size Dev size Test size Reasoning type Source Image-dependent Targeted affordance Publicly available
PaCo Qasemi et al. (2022a) 5,580 1,860 4,960 Preconditioned commonsense Crowd-sourced
WINOVENTI Do and Pavlick (2021) - - 4,352 Commonsense with exceptions Crowd-sourced
PVLIR Qasemi et al. (2023) 34,000 Preconditioned visual commonsense Other dataset
Normlense Han et al. (2023) - - 10,000 Defeasible visual commonsense Crowd-sourced
WinoViz Jin et al. (2024) - - 1,380 Reasoning object’s visual property Crowd-sourced
PROST Aroca-Ouellette et al. (2021) - - 18,736 Reasoning object’s physical property Other dataset
NEWTON Wang et al. (2023) - - 2,800 Reasoning object’s physical property Crowd-sourced
Persiani and Hellström (2019) 734,002 - 314,572 Object affordance without context Synthetic
Text2Afford (Ours) - - 35,520 (2368 * 15) Contextual object affordance Crowd-sourced
Table 1: Comparison of Text2Afford with other reasoning datasets. A larger version is in Appendix A.3 Table 9.
@@
As another example, for the sentence “an apple in the tree”, we should infer that the “apple” can be eaten, and is rollable. However we cannot roll an “apple logo”. In computer vision and robotics efforts, an accompanying image (or video) often provides necessary information about shape and physical properties of the entity, which can be used to predict affordances Zhu et al. (2014). However such information is absent in NLP tasks. To capture this nuance, we annotate crowdsourced text intended for other tasks (such as NLI) with the objects and affordances. We use 15 affordance classes from Zhu et al. (2014). Through extensive pilot studies, we train a set of annotators using the toloka.ai platform. We choose $25$ highly-skilled annotators who annotated a total of $2368$ sentence-object pairs with 15 affordance classes, on a 0-3 Likert-like scale. For each sentence-object pair and each affordance class we ensure annotations from three annotators to enable majority votings. We name this novel dataset Text2Afford. We use the dataset for zero-shot evaluations of small LMs, open-source LLMs and also some VLMs by forming different task setups. Figure 1 presents an example from Text2Afford and the derived tasks (detailed in Section 4). We evaluate the effect of few-shot fine-tuning on few PTLMs and VLM. Our contributions can be summarized as follows.
@@
$\bullet$ We curate a novel large scale crowdsourced text to affordance dataset – Text2Afford, consisting of 35,520 test data points (2368 sentence-object pairs with 15 unique affordance classes per pair). We ensure at least three annotations for each sentence-object pair for each class.
@@
$\bullet$ Using Text2Afford, we perform zero-shot evaluation of several state-of-the-art PTLMs along with a few VLMs in different settings to identify the extent to which they gain the knowledge of affordance during pretraining. We further ensemble the VLM and the PTLM predictions to examine whether pre-training with images can enrich affordance prediction from text. Overall, we observe that the SOTA LLMs face difficulties predicting contextual object affordances solely from text (accuracy $<55\%$) and the performance gets slightly enhanced when using powerful VLMs in presence of synthetic images.
@@
$\bullet$ We also fine-tune few PTLMs on a small subset of our data as well as on some commonsense reasoning tasks to understand how quickly the affordance knowledge get scaled up and how far the affordances are related to commonsense knowledge. In addition, we examine the in-context learning (ICL) ability of few of the SOTA generative LLMs and VLMs in affordance prediction task. We find that the pre-trained encoder based models gain some knowledge about object affordance during fine-tuning using the commonsense reasoning dataset.
$\bullet$ Additionally through finetuning on our dataset, we show that knowledge of affordance can improve model’s physical reasoning capability.
@@ 2 Related work
Reasoning about object affordances. Object affordances has been extensively studied in Computer Vision and Robotics (Sun et al. (2014); Zhu et al. (2014)). Recent methods employ deep learning approaches to detect object affordance. Nguyen et al. (2017) applies an object detector, CNN and dense conditional random fields to detect object affordance from RGB images. Persiani and Hellström (2019) extracts object-action pairs from web corpora using semantic role labelling. In contrast, we propose a crowd-sourced text only affordance dataset to audit the strength of SOTA LLMs and VLMs to reason about contextualized object affordance.
@@
Probing methods. Talmor et al. (2020) utilizes probing and employs Multi-choice MLM (Masked Language Modelling) and Multi-choice QA (Question Answering) setup to capture reasoning capabilities of pre-trained Language Models. Yang et al. (2022) examines zero-shot prediction performances on different tasks by LLM through novel visual imagination. Aroca-Ouellette et al. (2021) highlights the shortcomings of state-of-the-art pre-trained models in physical reasoning, with a further performance decline observed when incorporating option shuffling and superlatives in reasoning questions. Liu et al. (2022) proposes a novel spatial commonsense probing framework to investigate object scales and positional relationship knowledge in text-based pre-trained models and models with visual signals.
Reasoning tasks and dataset.  Reasoning about object affordance require a sort of commonsense reasoning. A series of works (Singh et al. (2021), et al. (2023), Bisk et al. (2019), Huang et al. (2019), Talmor et al. (2019), Talmor et al. (2021), Zellers et al. (2018)) study the text based commonsense knowledge of language models. Dataset such as $\delta$-NLI Rudinger et al. (2020) focuses on defeasible inference of commonsense knowledge; PaCo Qasemi et al. (2022a) and PInKS Qasemi et al. (2022b) deal with preconditioned commonsense inference of language models. PVLIR Qasemi et al. (2023), Normlense Han et al. (2023) use images as precoditions to reason about defeasible commonsense norms. However, none of these specifically focus on reasoning of object affordance. Wang et al. (2023) proposes a benchmark of object-attribute pairs plus a diverse set of questions to reason object’s physical properties. Aroca-Ouellette et al. (2021) tackles physical and affordance reasoning from an object-centric approach. Persiani and Hellström (2019) attempts to extract common object-action pairs from web corpora. In Table 1, we demonstrate the comparison of Text2Afford with other datasets which perform different kind of reasoning tasks. We emphasize that, Text2Afford is the largest crowdsourced publicly available text based contextualized affordance dataset with a test size of 35,520 (2368 sentence-object pairs and 15 affordance classes).
Present work. Although a substantial number of work study the reasoning capabilities of language models and propose commonsense reasoning datasets, however, none of these work concentrate specifically on evaluating the knowledge of affordance and contextual affordance prediction solely from text. To bridge this gap, we present a reliable crowdsourced test dataset for identifying the contextualized affordance prediction capability of LLMs as well as VLMs. Our results show that the advanced large language models fail to understand an object’s physical properties aka the affordances from texts, and there is significant room for improvement which may further motivate researchers to explore models that explicitly learns to ground objects in text to predict its physical properties and affordances.
  Agreement category Affordance classes Objects Object-affordance pair
High agreement (>0.75) Row, Feed, Ride, Fix the horse, striped white shirts, a brown paper sack, Chinese lanterns, Adrin’s sword, The movie breakfast-Feed, a horse-Watch, crops-Fix, sports-Grasp, sports-Lift, sports-Push, the phone-Feed, football-Ride
Medium agreement (>0.4 & <0.75) Throw, PourFrom, WriteWith, LookThrough, Lift, Grasp, Play, Push A red flag, An arrow, Art, Automatic weapons, Babies, Black-and-white TV computers-WriteWith, cats-Feed, football-Play, book-WriteWith, the door-Push
Low agreement (<0.4) Watch, SitOn, TypeOn Brandy from Spain, stone circles, iron, batteries, his fist, historical artifact, gift, olive oil, outdoor tables, bumper sticker on a car weapon-Push, The table-Lift, boat-Fix, paintings-LookThrough, cats-Throw
Table 2: Agreement based on difficulty in disambiguating different affordance classes, objects and object-affordance pairs.
3 Text2Afford dataset construction
@@
We select $20,000$ sentences from a crowdsourced English dataset (XNLI English) Conneau et al. (2018) [^2] We choose XNLI as a source to facilitate multilingual extensions of our dataset. and extract the noun phrases using the Stanford CoreNLP tool. As we restrict to the affordances that humans can directly perform, we filter the phrases which do not represent a tangible object (using ConceptNet). We manually filter out objects that cannot be acted upon directly by humans (such as school, building). After this preprocessing, we obtain a set of sentence-object pairs ($\langle x_{i},o_{i}\rangle$), where the sentence acts as the context for the corresponding object. Each sentence on average has 2-3 such objects. We use the 15 predefined affordance classes from Zhu et al. (2014) to label each sentence-object pair for annotation.
We utilize the Toloka platform [^3] https://toloka.ai/ for conducting the data annotation. We design an interface on this platform, containing clear instructions and examples for annotating the data. We conduct two rounds of pilot studies along with additional AMA (Ask Me Anything) sessions to analyze the subjective understanding of the annotators and, thereby, only select the high quality, serious annotators. A total of 114 annotators participated in the pilot study, and out of that we finally engage 25 skilled annotators to annotate a total of 2,368 sentence-object pairs each containing 15 affordance classes. Each datapoint (i.e., sentence-object pair along with an affordance class) has been annotated by three different annotators. We provide the details of the pilot studies & annotator training in Appendix A.1. By evaluating the complexity of the task for the annotators from the pilot studies, we intentionally consider a relatively small number of datapoints at a point for the annotation. This leads us to a total of 10 phases to complete the final annotation. We carefully reviewed each annotation and provided feedback with guidance in case of mistakes. For instance, annotators initially got confused with the affordance ‘Watch’ as human can watch any visual objects. In another instance, some annotators asked whether ‘Throw’ can be valid affordance for the object ‘Kittens’ as humans can perform ‘Lift’, ‘Throw’ to the object ‘Kitten’. We discussed these types of ambiguities with the annotators after each phase. Throughout each of the data annotation phases, we put scrupulous attention to quality control, including iterative annotation refinement, and manual evaluation. The overall statistics for this currently constructed dataset – Text2Afford is in Table 3. The Text2Afford dataset consists of $2368$ sentence-object pairs having $\sim 100k$ annotations (2368 $\times$ 15 $\times$ 3). For further details of the dataset construction, and our method of handling ambiguous scenarios, we refer the reader to Appendix A.1.
@@
Table 3: Text2Afford dataset statistics. # of instances annotated: (# of <s-o> pairs) * (# of classes) * (# of annotations per class).
@@
Text2Afford data exploration. We observe that classes such as ‘Grasp’, ‘Lift’, ‘Throw’, ‘Push’, and ‘Watch’ are the most common affordances for the objects present in the dataset (see Figure 2). Most frequent objects and their corresponding agreement scores are shown in Appendix A.10 Fig. 8. We observe, agreement scores are fairly uniform ($0.5$-$0.6$) for frequent objects, with high agreement for some frequent objects ($0.8$ for “the movie”). In Figure 9 (see Appendix A.9), we also see that ‘Grasp’, ‘Lift’, and ‘Throw’ are highly correlated classes. There is similar positive correlation between the class ‘SitOn’ and ‘Ride’, and some correlation between ‘Watch’ and ‘LookThrough’.
In Table 2, we list down the affordance classes based on the annotator agreement score, and divide it into three categories to understand which of the affordance classes pose the most and least difficulties for the human annotators. We observe that the classes - ‘Watch’, ‘SitOn’, and ‘TypeOn’ are the most difficult to disambiguate. Further, to explore the difficulty of understanding contextual object affordance, we employ three naïve annotators to annotate some samples of the Text2Afford, and we observe that on an average in 88% cases the humans are able to predict affordance correctly, and in some cases the context introducing inherent difficulty for predicting affordance. Details of the study is provided in Appendix A.2.
4 Task description
Our first objective is to audit the strength of Large Language Models in identifying the pre-defined affordance classes of objects from text in zero-shot settings. Given a textual context, and the object, the task is to predict whether a particular affordance class is applicable to that object conditioned on the context. We majorly leverage 4 types of task setup for the experiments. For the encoder based models (e.g., RoBERTa, BERT) we choose Masked Language Modelling (MLM) and Natural Language Inference (NLI) based setup, and for the generative models we adopt 2 types of probing setup (text only, text+image) to formalize the task. Table 4 demonstrates different types of tasks that we engage for conducting the experiments from the Text2Afford dataset.
  Model architecture Tasks Input instance Output
Encoder based MLM All the women in India wear bangles [SEP_TOKEN] bangles can be used for [MASK_TOKEN] by human Probabilities of each affordance classes as the [MASK_TOKEN]
NLI premise: All the women in India wear bangles Entailment scores for each affordance classes
hypothesis: bangles can be used for <Affordance> by human
[…]
Table 4: Overview of the tasks using Text2Afford. For detailed prompting see Appendix 10.